OpenStudy (anonymous):

I need a question checked, would it be A and B Which literal equations are equivalent to k=f/x? Choose all answers that are correct. A.F=kx B. x=kf C.F/kx D. x=fx

OpenStudy (anonymous):

D is supposed to be x= f/x

OpenStudy (anonymous):

And C is F=k/x I got that mixed up

OpenStudy (tkhunny):

Have you considered solving all of them for k and seeing which are the same?

OpenStudy (anonymous):

Yeah about that, I don't know how to... @tkhunny

OpenStudy (tkhunny):

Not good at all. You must have missed a unit. A. f = kx Solve for k. x is attached to k via multiplication. Use division to get it off. Divide by x f/x = kx/x = k ==> f/x = k Is that equivalent? You do the rest.

OpenStudy (anonymous):

Well yeah it is, it's just switched